# Towards Unbiased and Robust Spatio-Temporal Scene Graph Generation and Anticipation

### v1

@@ 4 Technical Approach
We propose ImparTail, a framework for unbiased Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA). In ImparTail, we adopt an object-centric relationship representation processing pipeline as illustrated in Fig 2. Thus, it can be easily extended to de-bias any VidSGG or SGA method built on an object-centric relationship representation processing pipeline [^2] Most VidSGG [7, 10, 30, 21] and SGA [35] approaches share a similar object-centric relationship representation processing pipeline. We further note that recent SOTA VidSGG methods still rely on the object-centric pipelines introduced by STTran [7] and DSGDetr [10].. ImparTail comprises four key components. (I) Object Representation Processing Unit (ORPU): This module generates representations for objects detected in video frames. (II) Spatio-Temporal Context Processing Unit (STPU): That constructs object-centric relationship representations tailored for two tasks: observed relationships for VidSGG and anticipated relationships for SGA. (III) Relationship Predicate Decoders: These decoders predict the relationship predicate labels corresponding to the observed or anticipated relationship representations [^3] Note that ORPU, STPU and Predicate Decoders can be adapted from any VidSGG or SGA methods utilizing an object-centric pipeline.. (IV) Curriculum-Guided Masked Loss: We implement a curriculum-based masking approach that selectively omits the contribution of specific relationship predicate labels during each training iteration. Thus, this method allows the model to progressively balance the class distribution by concentrating on underrepresented classes.
The following sections provide a comprehensive overview of the key components and training methodologies of ImparTail. Specifically, In Sec. 4.1, we outline a detailed explanation of ORPU and STPU, including their respective inputs and outputs within a single training iteration. In Sec. 4.2, we provide a detailed description of the proposed Curriculum-Guided Mask Generation Strategy, and in Sec. 4.3, we present the final change imparted in the loss function for both VidSGG and SGA tasks.
Figure 3: (a) Robustness Evaluation Pipeline: We outline a methodology to evaluate the robustness of trained VidSGG and SGA models under input distribution shifts. Specifically, we introduce systematic corruptions into the frames of test videos, which are then fed as input to the trained models. (b) Corrupted Frames: We illustrate the resultant frames obtained by inducing various categories of corruptions.
@@ 4.1 ImparTail Pipeline
Both ORPU and STPU are designed to be compatible with any VidSGG or SGA method. As most existing approaches use similar processing pipelines, we outline the essential architectural components required for our framework to effectively de-bias their methods [^4] Current methods in the literature typically incorporate at least one additional component beyond those specified here. Our experiments directly integrated the ORPU and STPU units from the selected methods.
(I) Object Representation Processing Unit (ORPU): We first extract visual features of observed objects using a pre-trained object detector. Specifically, we extract visual features ($\textstyle\{\mathbf{v}^{t}_{i}\}_{i=1}^{N(t)}$), bounding boxes ($\textstyle\{\mathbf{b}^{t}_{i}\}_{i=1}^{N(t)}$), and object category distributions ($\textstyle\{\hat{\mathbf{c}}^{t}_{i}\}_{i=1}^{N(t)}$), for object proposals $\textstyle\{o^{t}_{i}\}_{i=1}^{N(t)}$ in the observed frames. We use a transformer encoder [43] to integrate temporal information and generate consistent object representations. It is given by:
@@
(II) Spatio-Temporal Context Processing Unit (STPU): To learn spatio-temporal context-aware relationship representations, we first construct representations for interacting objects, then refine spatial context with an encoder, and finally use a transformer encoder to integrate spatio-temporal context. Specifically, we define the relationship representation for each observed frame $\textstyle I^{t}$ as $\textstyle\mathbf{z}_{ij}^{t}$, where $\textstyle\mathbf{Z}^{t}$ is formed by stacking all relationship features $\textstyle\{\mathbf{z}_{ij}^{t}\}_{ij}$ corresponding to the objects within the frame. This stacked representation $\textstyle\mathbf{Z}^{t}$ is then passed through the transformer encoder to produce spatial-context-aware representations. Next, we construct the matrix $\textstyle\mathbf{Z}_{ij}$ by stacking the relationship representations $\textstyle\{\mathbf{z}_{ij}^{t}\}_{t=1}^{T}$ over all observed frames. This matrix $\textstyle\mathbf{Z}_{ij}$ is then passed through an encoder, which aggregates temporal information across the stacked relationship representations:
@@
(III) Relationship Predicate Decoders: We utilize a two-layer multi-layer perceptron (MLP) to decode the output relationship representations from STPU. We note that for SGA task, following SceneSayer [35], we employ two relationship predicate decoders. One for the observed relationship representations and the other for the anticipated representations. They can be described as follows:
@@
4.2 Curriculum-Guided Mask Generation
@@
Devising curriculum learning approaches for any problem necessitates precise, problem-specific definitions of three key components [28]: (a) the ability to rank tasks from easy to hard, (b) the capability to determine mastery over a specific task, and (c) a procedure that periodically integrates easier tasks with more challenging ones. In this vein, ImparTail employs a simple yet significantly effective curriculum learning method to generate masks for relationship predicate labels during training VidSGG and SGA models.
@@
Banking on the transformers’ prowess in sequential processing tasks, we design the curriculum-guided mask generation process for VidSGG and SGA tasks using simple heuristics as follows: (a) Ranking: We assume that more data associated with a particular relationship predicate label indicates a better ability to predict that class. Therefore, we rank the hardness of relationship predicate learning based on the quantity of data available for them. (b) Mastery: We employ recall as the metric to assess mastery over predicting a specific relationship predicate class, while mean recall evaluates overall performance across all classes. (c) Progressive Mixing: We implement a distributional strategy that progressively balances the class distribution, gradually moving toward an equilibrium where all classes are uniformly represented. The above discussion yields Alg 1, where, as described, in each iteration, we progressively increase the number of masks over dominant relationship predicate labels, voiding their contribution.
4.3 Masked Loss a.k.a Partial SGD
During training, at each epoch $\textstyle e$, as described above, curriculum-guided $\textstyle\operatorname{MaskGenerator}$ provides us with a set of masks $\textstyle\mathcal{M}^{(e)}$ for all the video segments in the dataset $\textstyle\mathcal{D}$. Now, let $\textstyle\mathcal{L}_{p^{t}_{ij}}$ represent any loss defined over a single relationship predicate. Upon application of the mask, the resultant masked loss $\textstyle\mathscr{L}_{p^{t}_{ij}}$ is given by:
$\mathscr{L}_{p^{t}_{ij}}=\mathbf{m}^{t}_{ij}*\mathcal{L}_{p^{t}_{ij}}=\begin{cases}\mathcal{L}_{p^{t}_{ij}}&\text{if }\mathbf{m}^{t}_{ij}=1\\ 0&\text{if }\mathbf{m}^{t}_{ij}=0\end{cases}$ (4)
@@
Input: Epoch: $\textstyle e$, Sampling ratio: $\textstyle\mathcal{S}$, Dataset Annotations: $\textstyle\mathcal{D}$
Output: Ground Truth Annotation Masks: $\textstyle\mathcal{M}$
@@
2Compute total annotations $\textstyle N=\sum_{v,f}|\mathcal{D}[v][f]|$
3
4*** Determine Target Counts ***
@@
9*** For an equally weighted multinomial distribution ***
@@
14*** Randomly sample instances of relationship predicates in the dataset based on the target counts for each relationship predicate and construct the filtered dataset ***
@@
16Construct filtered dataset $\textstyle\mathcal{F}$
@@
18*** Reconstruct Filtered Ground Truth Annotations ***
19 Initialize empty mask list $\textstyle\mathcal{M}$
20 foreach Video $\textstyle v$ in $\textstyle\mathcal{G}$ do
@@ 25                   Initialize object mask $\textstyle\mathcal{M}_{o}$
26                   Original relations $\textstyle\mathcal{R}_{o}$ from $\textstyle\mathcal{D}$
27                   Filtered relations $\textstyle\mathcal{F}_{o}$ from $\textstyle\mathcal{F}_{r}$
@@
29                         if $\textstyle rel\in\mathcal{F}_{o}[v][f]$ then
@@ 38            Add $\textstyle\mathcal{M}_{f}$ to $\textstyle\mathcal{M}_{v}$
39
40      Add $\textstyle\mathcal{M}_{v}$ to $\textstyle\mathcal{M}$
@@
VidSGG.
models trained using ImparTail, comprise the combination of (1) Object Classification Loss on the object representations and (2) Masked Predicate Classification Loss on the observed relationship representations as their training objective [^5] Object and predicate losses above can be adapted from any method..Thus, the resultant objective is:
$\underbrace{\mathcal{L}_{i}=\sum_{t=1}^{\bar{T}}\mathcal{L}_{i}^{t}}_{\text{(1)}};\underbrace{\mathscr{L}_{\text{gen}}=\sum_{t=1}^{\bar{T}}\mathscr{L}_{\text{gen}}^{t},\mathscr{L}_{\text{gen}}^{t}=\sum_{ij}\mathbf{m}^{t}_{ij}*\mathcal{L}_{p^{t}_{ij}}}_{\text{{Masked} Predicate Classification Loss (2)}}\quad\\ \mathcal{L}=\sum_{t=1}^{\bar{T}}\left(\lambda_{1}\mathscr{L}_{\text{gen}}^{t}+\lambda_{2}\sum_{i}\mathcal{L}_{i}^{t}\right)$
$\underbrace{\mathcal{L}_{i}=\sum_{t=1}^{\bar{T}}\mathcal{L}_{i}^{t},}_{\text{(1)}};\underbrace{\mathscr{L}_{\text{gen}}=\sum_{t=1}^{\bar{T}}\mathscr{L}_{\text{gen}}^{t},\mathscr{L}_{\text{gen}}^{t}=\sum_{ij}\mathbf{m}^{t}_{ij}*\mathcal{L}_{p^{t}_{ij}}}_{\text{{Masked} Observed Predicate Classification Loss (2)}}\\ \underbrace{\mathscr{L}_{\text{ant}}^{(1:T)}=\sum_{t=T+1}^{\min(T+H,\bar{T})}\mathscr{L}_{\text{ant}}^{t},\quad\mathscr{L}_{\text{ant}}^{t}=\sum_{ij}\mathbf{m}^{t}_{ij}*\mathcal{L}_{p^{t}_{ij}}}_{\text{{Masked} Anticipated Predicate Classification Loss (3)}}\\ \mathcal{L}=\sum_{t=1}^{\bar{T}}\left(\lambda_{1}\mathscr{L}_{\text{gen}}^{t}+\lambda_{2}\sum_{i}\mathcal{L}_{i}^{t}\right)+\\ \sum_{T=3}^{\bar{T}-1}\left(\lambda_{3}\mathscr{L}_{\text{ant}}^{(1:T)}+\lambda_{4}\mathcal{L}_{\text{boxes}}^{(1:T)}+\lambda_{5}\mathcal{L}_{\text{recon}}^{(1:T)}\right)$ (5)
SGA.
models trained using ImparTail, comprise a combination of masked losses over two types of relationship representations: (a) Observed Relationship Representations. and (b) Anticipated Relationship Representations. and other method specific losses. Thus, the objective when ImparTail is employed to SceneSayer [35] includes (1) Object Classification Loss, (2) Masked Predicate Classification Loss over observed relationship representations, (3) Masked Predicate Classification Loss over anticipated relationship representations, (4) Bounding Box Regression Loss and (5) Reconstruction Loss and is given in Eq 5.

### v2

@@ 4 Technical Approach
In ImparTail, we adopt an object-centric relationship representation processing pipeline as illustrated in Fig 3. Thus, it can be easily extended to de-bias any VidSGG/SGA method [^4] Note: ImparTail does not train models for both tasks simultaneously. Instead, it adapts existing approaches for VidSGG/SGA separately and learns the corresponding VidSGG/SGA unbiased estimators. built on an object-centric relationship representation processing pipeline [^5] Most VidSGG [7, 11, 32, 23] and SGA [37] approaches share a similar object-centric relationship representation processing pipeline. We further note that recent SOTA VidSGG methods still rely on the object-centric pipelines introduced by STTran [7] and DSGDetr [11].. ImparTail comprises four key components. (I) Object Representation Processing Unit (ORPU): This module generates representations for objects detected in video frames. (II) Spatio-Temporal Context Processing Unit (STPU): That constructs object-centric relationship representations tailored for two tasks: observed relationships for VidSGG and anticipated relationships for SGA. (III) Relationship Predicate Decoders: These decoders predict the relationship predicate labels corresponding to the observed or anticipated relationship representations [^6] Note: As ImparTail adapts existing VidSGG/SGA methods, we re-use the feature extraction strategy, ORPU, STPU and Predicate decoder blocks employed by corresponding object-centric VidSGG/SGA methods.. (IV) Masked Training with Partial Gradients: We propose an impartial training objective based on loss masking that curbs the influence of dominant head classes. These are the classes that are over-represented in the dataset and often lead to biased predictions. Our approach aims to reduce their influence while focusing on under-represented tail relationships. Furthermore, our curriculum-driven mask generation method enables the model to iteratively refine its bias mitigation strategy over time, resulting in a balanced performance.
The following sections provide a comprehensive overview of the key components and training methodologies of ImparTail. Specifically, In Sec. 4.1, we outline a detailed explanation of ORPU and STPU, including their respective inputs and outputs within a single training iteration. In Sec. 4.2, we provide a detailed description of the proposed Masked Training with Partial Gradients (refer Appendix Sec.4 for details on the specific changes employed to STSG methods).
@@ 4.1 ImparTail Pipeline
Given that our approach builds upon existing object-centric VidSGG/SGA methods—most of which share similar representation processing pipelines—we detail the essential architectural components necessary for our framework to effectively adapt a method and learn an unbiased estimator [^7] Current methods in the literature typically incorporate at least one additional component beyond those specified here. ImparTail directly integrates the ORPU, STPU units from them ignoring additional components.
(I) Object Representation Processing Unit (ORPU): extracts visual features ($\textstyle\{\mathbf{v}^{t}_{i}\}_{i=1}^{N(t)}$), bounding boxes ($\textstyle\{\mathbf{b}^{t}_{i}\}_{i=1}^{N(t)}$), and object category distributions ($\textstyle\{\hat{\mathbf{c}}^{t}_{i}\}_{i=1}^{N(t)}$), for object proposals $\textstyle\{o^{t}_{i}\}_{i=1}^{N(t)}$ in the observed frames utilizing a pre-trained object detector proposed by specific VidSGG/SGA method employed for adaptation. It then employs a transformer encoder [46] to integrate temporal information and generate temporally consistent object representations (see Eq. 1):
@@
(II) Spatio-Temporal Context Processing Unit (STPU): learns spatio-temporal context-aware relationship representations, it first constructs representations for interacting objects, then refines spatial context with an encoder, and finally use a transformer encoder to integrate spatio-temporal context information. Specifically, let the relationship representation for each observed frame $\textstyle I^{t}$ as $\textstyle\mathbf{z}_{ij}^{t}$, where $\textstyle\mathbf{Z}^{t}$ is formed by stacking all relationship features $\textstyle\{\mathbf{z}_{ij}^{t}\}_{ij}$ corresponding to the objects within the frame. $\textstyle\mathbf{Z}^{t}$ is then passed through the transformer encoder to produce spatial-context-aware representations. Next, it constructs the matrix $\textstyle\mathbf{Z}_{ij}$ by stacking the relationship representations $\textstyle\{\mathbf{z}_{ij}^{t}\}_{t=1}^{T}$ over all observed frames. This matrix $\textstyle\mathbf{Z}_{ij}$ is then passed through a temporal encoder, which aggregates temporal information across the representations (refer LDPU [37] for SGA task):
@@
(III) Relationship Predicate Decoders: Current VidSGG methods utilize a two-layer MLPs to decode the output relationship representations from STPU described as follows:
$\hat{\mathbf{p}}^{t}_{ij}=\operatorname{PredClassifier}\left(\mathbf{z}^{t}_{ij}\right),\forall t\in[1,\bar{T}]$ (3)
We note that for SGA task, following SceneSayer [37], we employ two relationship predicate decoders. One for the observed relationship representations and the other for the anticipated representations described as follows:
@@
4.2 Masked Training with Partial Gradients
Our approach stems from the observation that instead of re-weighting training data points which, in our case, entails uniformly adjusting the loss contributed by all relationship predicates derived from observed object pairs across all frames of video), selectively masking the loss contributed by head relationship classes helps learn unbiased estimators.
4.2.1 Masked Loss
During training, for epoch $\textstyle e$ and mask-ratio $\textstyle\mathscr{R}_{m}$, Mask Generator generates a set of relationship predicate label masks $\textstyle\mathcal{M}^{(e)}=\{\{\mathbf{m}^{t}_{ijk}\}_{ijk}\}_{t}$ where $\textstyle\mathbf{m}^{t}_{ijk}\in\{0,1\}$. Let $\textstyle\mathcal{L}_{p^{t}_{ijk}}$ represent any loss defined over a relationship predicate888(a) Two interacting objects (i,j) at time (t) can have multiple relationship predicates $\textstyle\{p^{t}_{ijk}\}_{k}$ that describe the nature of the interaction; (b) We re-use the predicate loss $\textstyle\mathcal{L}_{p^{t}_{ijk}}$ from the method employed for adaptation.. A mask initialized i.e. setting $\textstyle\mathbf{m}^{t}_{ijk}=1$ voids the contribution of loss $\textstyle\mathcal{L}_{p^{t}_{ijk}}$ and results in masked loss $\textstyle\mathscr{L}_{p^{t}_{ij}}$ given by:
$\mathscr{L}_{p^{t}_{ijk}}=(1-\mathbf{m}^{t}_{ijk})*\mathcal{L}_{p^{t}_{ijk}}=\begin{cases}\mathcal{L}_{p^{t}_{ijk}}&\text{if }\mathbf{m}^{t}_{ijk}=0\\ 0&\text{if }\mathbf{m}^{t}_{ijk}=1\end{cases}$ (5)
VidSGG. models trained using ImparTail, comprise the combination of (1) Object Classification Loss on the object representations and (2) Masked Predicate Classification Loss on the observed relationship representations as their training objective.The resultant objective is given in Eq 6:
$\underbrace{\mathcal{L}_{i}=\sum_{t=1}^{\bar{T}}\mathcal{L}_{i}^{t}}_{\text{(1)}};\quad\underbrace{\mathscr{L}=\sum_{t=1}^{\bar{T}}\mathscr{L}^{t},\mathscr{L}^{t}=\sum_{ijk}\mathscr{L}_{p^{t}_{ijk}}}_{\text{{Masked} Predicate Classification Loss (2)}}\quad\\ \mathcal{L}=\sum_{t=1}^{\bar{T}}\left(\lambda_{1}\mathscr{L}^{t}+\lambda_{2}\sum_{i}\mathcal{L}_{i}^{t}\right)$ (6)
SGA. models trained using ImparTail, comprise a combination of masked losses over two types of relationship representations: (a) Observed Relationship Representations. and (b) Anticipated Relationship Representations. and other method specific losses. Thus, the objective when ImparTail is employed to SceneSayer [37] includes (1) Object Classification Loss, (2) Masked Predicate Classification Loss over observed relationship representations, (3) Masked Predicate Classification Loss over anticipated relationship representations, (4) Bounding Box Regression Loss and (5) Reconstruction Loss as shown in Eq 7.
$\underbrace{\mathcal{L}_{i}=\sum_{t=1}^{\bar{T}}\mathcal{L}_{i}^{t}}_{\text{(1)}};\quad\underbrace{\mathscr{L}_{\text{gen}}=\sum_{t=1}^{\bar{T}}\mathscr{L}_{\text{gen}}^{t},\mathscr{L}_{\text{gen}}^{t}=\sum_{ijk}\mathscr{L}_{p^{t}_{ijk}}}_{\text{{Masked} Observed Predicate Classification Loss (2)}}\\ \underbrace{\mathscr{L}_{\text{ant}}^{(1:T)}=\sum_{t=T+1}^{\min(T+H,\bar{T})}\mathscr{L}_{\text{ant}}^{t},\quad\mathscr{L}_{\text{ant}}^{t}=\sum_{ijk}\mathscr{L}_{p^{t}_{ijk}}}_{\text{{Masked} Anticipated Predicate Classification Loss (3)}}\\ \mathcal{L}=\sum_{t=1}^{\bar{T}}\left(\lambda_{1}\mathscr{L}_{\text{gen}}^{t}+\lambda_{2}\sum_{i}\mathcal{L}_{i}^{t}\right)+\\ \sum_{T=3}^{\bar{T}-1}\left(\lambda_{3}\mathscr{L}_{\text{ant}}^{(1:T)}+\lambda_{4}\mathcal{L}_{\text{boxes}}^{(1:T)}+\lambda_{5}\mathcal{L}_{\text{recon}}^{(1:T)}\right)$ (7)
4.2.2 Mask Generation
@@
We observed that by tuning the mask-ratio ($\textstyle\mathscr{R}_{m}$) (i.e., the fraction of annotations masked), ImparTail can learn estimators along a spectrum, from those exhibiting bias to those that are unbiased. We propose curriculum-guided mask generation to enable models to progressively refine their bias mitigation strategy. Devising curriculum learning approaches for any problem necessitates precise, problem-specific definitions of three key components [31]: (a) the ability to rank tasks from easy to hard, (b) the capability to determine mastery over a specific task, and (c) a procedure that periodically integrates easier tasks with more challenging ones. Thus, ImparTail employs a simple yet effective curriculum learning method to generate masks for relationship predicate labels to train VidSGG/SGA models.
@@
Banking on the transformers’ prowess in sequential processing tasks, we design the curriculum-guided mask generation process for VidSGG and SGA tasks using simple heuristics as follows: (a) Ranking: We assume that more data associated with a particular relationship predicate label indicates a better ability to predict that class. Therefore, we rank the hardness of relationship predicate learning based on the quantity of data available for them. (b) Mastery: We employ recall as the metric to assess mastery over predicting a specific relationship predicate class, while mean recall evaluates overall performance across all classes. (c) Progressive Mixing: We propose a simple distributional strategy that progressively balances the class distribution, gradually moving toward an equilibrium where all classes are uniformly represented. Simply put, as training progresses, we stochastically mask the loss of head classes (easier examples), allowing the model to focus on learning about tail classes (harder examples) and the number of masks increases each epoch. The above discussion yields Alg 3.
@@
Input: Epoch: $\textstyle e$, Sampling Ratio: $\textstyle\mathscr{R}_{s}$, Dataset Annotations: $\textstyle\mathcal{D}$, Total predicate labels: $\textstyle N$, Total predicates: $\textstyle\mathcal{P}$, Videos: $\textstyle\mathscr{V}$
Output: Masks: $\textstyle\mathcal{M}^{(e)}$
@@
2*** Determine Target Counts [can also be a fixed input] ***
[…]
4
@@
9*** Equally weighted distribution [can also be learnt] ***
@@
14*** Randomly sample instances of relationships in the dataset based on the target counts and construct filtered dataset ***
@@
16Construct filtered dataset $\textstyle\mathcal{F}$** See Appendix for details**
@@
18*** Construct masks such that any relation present in the filtered dataset is unmasked, while all others are masked. ***
19 Initialize empty mask list $\textstyle\mathcal{M}^{(e)}$
20 foreach Video $\textstyle v$ in $\textstyle\mathscr{V}$ do
@@ 25                   Initialize object mask $\textstyle\mathcal{M}_{o}$
26                   Original relations $\textstyle\mathcal{R}_{o}$ from $\textstyle\mathcal{D}[v][f]$
@@
28                         if $\textstyle rel\in\mathcal{F}[v][f]$ then
29                               Set mask value $\textstyle\mathcal{M}_{o}[rel]=0$
30
31                        else
@@ 37            Add $\textstyle\mathcal{M}_{f}$ to $\textstyle\mathcal{M}_{v}$
38
39      Add $\textstyle\mathcal{M}_{v}$ to $\textstyle\mathcal{M}^{(e)}$
40
@@ 4.3 Robustness Evaluation
Figure 4: (a) Robustness Evaluation Pipeline: We present a methodology to assess the robustness of trained VidSGG and SGA models when faced with input distribution shifts. Specifically, we systematically introduce corruptions to the frames of test videos, which are then used as inputs for the trained models. (b) Corrupted Frames: We illustrate the frames obtained by inducing various categories of corruptions (see Appendix Sec.4 for details).
5 Experiments
We apply ImparTail to generate and anticipate STSGs on the Action Genome dataset [19]. We prepare the dataset by following the pre-processing steps described in [7, 37].
In the following sections, we will assess the performance of ImparTail on four tasks, namely: (a) VidSGG, (b) SGA, (c) RobustVidSGG and (d) RobustSGA. For each task, we first select previously published methods, train and evaluate them as per the instructions, then adapt them to ImparTail framework, re-train and evaluate them. For VidSGG task, we selected STTran [7] and DSGDetr [11] as our baselines and for the SGA task, we selected STTran+ [37], STTran++ [37], DSGDetr+ [37], DSGDetr++ [37], SceneSayerODE [37], SceneSayerSDE [37] as our baselines.